# On the Three Demons in Causality in Finance: Time Resolution, Nonstationarity, and Latent Factors

### 28 Dec 2023 v1

@@ 4.3 Causal Discovery in the Presence of Latent Variables
Basically, Theorem 7 says that in certain conditions, we can build a unique connection between rank deficiency and atomic covers. Therefore, we can identify atomic covers by first ensuring the required conditions and then searching for combinations of $\mathcal{C}$ and $\mathcal{X}$ that induce rank deficiency. Next we briefly introduce how to design an algorithm that strives to fulfill these conditions, in order to cash out the rank-deficient property for identifying atomic covers in a graph, and consequently, identifying the whole causal structure over both latent and observed variables. More details can be found in our manuscript Dong et al. (2023).
Causal Discovery Algorithm. Our rank-based latent causal discovery consists of three phases.
Phase 1: Finding CI skeleton. In this phase we just follow the classical PC algorithm (Spirtes et al., 2000) to find the CI skeleton by using CI tests.
@@ 5 Experiments
(a) Change point detection for SP100 from 2008 to 2013. The detected change point in 2008 can be explained as Global financial crisis and subprime mortgage crisis starting from 2007. The detected change point in 2009 can be explained as European debt crisis starting from 2009.
@@ 5.1 Baysian Change Point Detection
To utilize our rank-based latent causal discovery method we need to assume that our data is stationary; otherwise the nonstationary would induce rank deficiency constraints that cannot correctly reflect the underlying causal structure. Therefore, we first employ Baysian change point detection Fearnhead (2006); Adams and MacKay (2007); Xuan and Murphy (2007) to identify the changing points in the time series data.
@@
Our results are shown in Figure 3, where most of the change points detected can relate to real-life financial crises or events. Specifically, the change point in 2008 can be explained as Global financial crisis and subprime mortgage crisis starting from 2007. The change point in 2009 can be explained as European debt crisis starting from 2009. The detected change point in 2016 could relate to Chinese stock market crash in 2016. The change point in 2018 could relate to cryptocurrency meltdown. The change point in 2020 likely relates to Covid-19. The change point in 2022 could relate to the Russian-Ukraine war.
@@
The detected change points in 2018 and 2020 are significant and are well related to real-life events. Thus we split the SP100 data using these two change points and have three time series: (i) SP100 from Jan 2017 to Dec 2017, (ii) SP100 from Jan 2019 to Dec 2019, and (iii) SP100 from Jan 2021 to Dec 2021. We will refer to them as SP100-2017, SP100-2019, and SP100-2021 respectively in the following.
@@
In the next section, we will show the causal discovery results using these three time series.
5.2 Rank-based Latent Causal Discovery
@@
Figure 4: Discovered causal graph using a subset of SP100 stocks for 2017, 2019, and 2021, respectively. We can see that these three graphs have many things in common, and yet a big difference between 2019 and 2021 is that there are many edges between travel-related stocks and pharmaceutical stocks after Covid-19, while before Covid-19 it is not the case.
We apply our rank-based latent causal discovery method to Sp100-2017, Sp100-2019, and Sp100-2021 respectively. The result is shown in Figure 4. We can see that these three graphs have things in common. For example, there is a latent variable influencing all the pharmaceutical companies and that detected substructure is consistent among Sp-2017, Sp-2019, and Sp-2021. We also notice that for Sp-2017, we do not have latent variable for tech companies but for Sp-2019 and Sp-2021 we do.
Furthermore there is a big difference between 2019 and 2021, which can be well-explained by Covid-19. We can see that there are many edges between travel-related stocks (BKNG and BA) and pharmaceutical stocks (BMY, PFE, JNJ, MRK, GILD, ABBV, CVS, LLY, AMGN) in the result using Sp100-2021, while for Sp100-2017 and Sp100-2019 it is not the case. This is because when pharmaceutical companies announce successful vaccine trials or gain approval for a COVID-19 vaccine, investor confidence in the travel sector increases, boosting airline stocks. On the other hand, if there are setbacks in vaccine development or distribution, airline stocks may suffer due to anticipated continued travel restrictions.

### 12 Jan 2024 v2

@@ 4.2 Causal Discovery in the Presence of Latent Variables
Basically, Theorem 7 says that in certain conditions, we can build a unique connection between rank deficiency and atomic covers. Therefore, we can identify atomic covers by first ensuring the required conditions and then searching for combinations of $\mathcal{C}$ and $\mathcal{X}$ that induce rank deficiency. Next we briefly introduce how to design an algorithm that strives to fulfill these conditions, in order to cash out the rank-deficient property for identifying atomic covers in a graph, and consequently, identifying the whole causal structure over both latent and observed variables. More details can be found in our manuscript (Dong et al., 2023).
Rank-based Latent Causal Discovery Algorithm. Our rank-based latent causal discovery consists of three phases.
Phase 1: Finding CI skeleton. In this phase we just follow the classical PC algorithm (Spirtes et al., 2000) to find the CI skeleton by using CI tests. Then we will rely on the information from the CI skeleton to portion the CI skeleton into subgraphs that might contain latent variables. Then we take these subgraphs as input to Phase 2.
(a) Take $\mathcal{X}=\{\{\mathsf{X_{3}}\}\}$ and
$\mathcal{C}=\{\{\mathsf{X_{8}}\}\}$.
(b) Take $\mathcal{X}=\{\{\mathsf{X_{2}},\mathsf{X_{3}}\}\}$ and $\mathcal{C}=\{\{\mathsf{X_{7}}\},\{\mathsf{X_{8}}\}\}$.
(c) Take $\mathcal{X}=\{\{\mathsf{X_{2}}\}\}$ and
$\mathcal{C}=\{\{\mathsf{X_{4}}\},\{\mathsf{X_{5}}\},\{\mathsf{X_{6}}\}\}$.
(d) Take $\mathcal{X}=\{\}$ and $\mathcal{C}=\{\{\mathsf{X_{1}}\},\{\mathsf{L_{2}},\mathsf{X_{2}}\},\{\mathsf{X_{3}}\}\}$
Figure 2: An illustrative example of the key process of our rank-based latent causal discovery algorithm.
@@
Phase 2 is the key procedure that identifies atomic covers and thus latent variables. Below we give an example with Figure 2 to illustrate the procedure.
Example 3
Consider the graph in Figure 2. We start with finding atomic covers with $k=1$, and we can find that $\{\mathsf{X_{3}}\}$ is a parent of $\{\mathsf{X_{8}}\}$, as in Figure 2(a). At this point, no more $k=1$ clusters can be found, so next we search for $k=2$ clusters. Then to identify collider $\{\mathsf{X_{7}}\}$, we only need to consider $\{\mathsf{X_{7}}\}$ and $\{\mathsf{X_{8}}\}$ together as the children of $\{\mathsf{X_{2}},\mathsf{X_{3}}\}$. After finding such a relationship, we arrive at Figure 2(b), and from now on the collider $\{\mathsf{X_{7}}\}$ will not induce unfavorable rank deficiency anymore (as it is recorded). The next step is to find the relation of $\{\mathsf{X_{4}}\}$,$\{\mathsf{X_{5}}\}$,$\{\mathsf{X_{6}}\}$ with $\{\mathsf{L_{2}},\mathsf{X_{2}}\}$, by taking $\mathcal{X}=\{\{\mathsf{X_{2}}\}\}$ and $\mathcal{C}=\{\{\mathsf{X_{4}}\},\{\mathsf{X_{5}}\}\}$ or $\{\{\mathsf{X_{4}}\},\{\mathsf{X_{6}}\}\}$ or $\{\{\mathsf{X_{5}}\},\{\mathsf{X_{6}}\}\}$, and thus conclude $\{\{\mathsf{X_{4}}\},\{\mathsf{X_{5}}\},\{\mathsf{X_{6}}\}\}$ as the pure children of $\{\mathsf{L_{2}},\mathsf{X_{2}}\}$, as in Fig 2(c). Finally we are able to find the relationship of $\{\mathsf{X_{1}}\}$,$\{\mathsf{L_{2}},\mathsf{X_{2}}\}$, $\{\mathsf{X_{3}}\}$ with $\{\mathsf{L_{1}}\}$, by taking $\mathcal{X}=\{\}$ and $\mathcal{C}$ as $\{\{\mathsf{X_{1}}\},\{\mathsf{L_{2}},\mathsf{X_{2}}\}\}$, $\{\{\mathsf{X_{1}}\},\{\mathsf{X_{3}}\}\}$, or $\{\{\mathsf{L_{2}},\mathsf{X_{2}}\},\{\mathsf{X_{3}}\}\}$, as in Figure 2(d).
@@
4.3 Determining All Directions and Estimating Causal Coefficients
Our rank-based latent causal discovery algorithm is able to output the Markove equivalence class of $\mathcal{O}_{\text{min}}(\mathcal{O}_{s}(\mathcal{G}))$. The skeleton of the graph is clear, and yet some directions of edges cannot be decided. Therefore, our next step is to determine all the rest undecided directions.
Determining Causal Directions. Here we aim at making use of Generalized Independent Noise (GIN) condition (Xie et al., 2020) to decide the directions, by assuming the noise terms $\epsilon_{\mathsf{V}_{i}}$ in Definition 1 are non-gaussian. The GIN condition is given as follows.
Condition 10 (GIN condition (Xie et al., 2020))
Assume an LLCM with non-gaussian additive noise. Let $\mathbf{Z}$ and $\mathbf{Y}$ be two observed random vectors. Define the surrogate for $\mathbf{Y}$ relative to $\mathbf{Z}$ as $E_{\mathbf{Y}||\mathbf{Z}}:=\omega^{T}\mathbf{Y}$, where $\omega$ satisfies $\omega^{T}\mathbb{E}[\mathbf{Y}\mathbf{Z}^{T}]=0$ and $\omega\neq 0$. We say that $(\mathbf{Z},\mathbf{Y})$ follows the GIN condition if and only if $E_{\mathbf{Y}||\mathbf{Z}}$ is independent from $\mathbf{Z}$.
It has been shown that, by making use of GIN conditions in specific ways, we can decide the directions of edges between latent variables Xie et al. (2020), summarized as follows.
Theorem 11 (Deciding Directions by GIN (Xie et al., 2020))
Assume an LLCM with non-gaussian additive noise. Let $\mathbf{V_{p}}$ and $\mathbf{V_{q}}$ be two atomic covers and there is no latent confounder behind $\mathbf{V_{p}}$ and $\mathbf{V_{q}}$, and $\mathbf{V_{p}}\cap\mathbf{V_{q}}=\varnothing$. Further suppose that $\mathbf{V_{p}}$ has $2|\mathbf{V_{p}}|$ pure children $\{\mathsf{P_{i}}\}_{1}^{2|\mathbf{V_{p}}|}$, and $\mathbf{V_{q}}$ has $2|\mathbf{V_{q}}|$ pure children $\{\mathsf{Q_{i}}\}_{1}^{2|\mathbf{V_{q}}|}$. If ($\mathbf{Z}=\{\mathsf{P_{i}}\}_{1}^{|\mathbf{V_{p}}|}$, $\mathbf{Y}=\{\mathsf{P_{i}}\}_{|\mathbf{V_{p}}|+1}^{2|\mathbf{V_{p}}|}\cup\{\mathsf{Q_{i}}\}_{1}^{|\mathbf{V_{q}}|}$) follows the GIN condition, then $\mathbf{V_{p}}\rightarrow\mathbf{V_{q}}$.
Theorem 11 is saying that even though we cannot directly observe latent variables, we can still decide the directions of edges between these latent variables by using observed variables as surrogates. Therefore, by first conducting our rank-based latent causal discovery algorithm and then checking directions following Theorem 11, we are able to discover the skeleton together with the directions of the underlying graph.
Estimating Causal Coefficients. Here we are interested in estimating the causal coefficients of each edge, given the causal structure and observational data. We provide our novel solution as follows. By the matrix form of Definition 1, i.e., $\mathbf{V}=A\mathbf{V}+\mathbf{\epsilon_{\mathbf{V}}}$, we have $\mathbf{V}=(\mathbf{I}-A)^{-1}\mathbf{\epsilon_{\mathbf{V}}}$, and $\Sigma_{\mathbf{V}}=(\mathbf{I}-A)^{-1}\Psi(\mathbf{I}-A)^{-T}$, where $\Psi$ is a diagonal matrix representing the covariance of $\epsilon_{V}$. We can take the submatrices from $\Sigma_{\mathbf{V}}$, and thus we have $\Sigma_{\mathbf{X}}$ and $\Sigma_{\mathbf{L}}$ parameterized by $A$ and $\Psi$. Similar to the typical factor analysis setting (Gorsuch, 2014), we assume that $\epsilon_{V}$ are all gaussian and thus $\mathbf{X}$ are jointly gaussian, and we find the causal coefficients $A$ by maximizing the log-likelihood of observational data:
$\arg\max_{A,\Psi}-\frac{n|\mathbf{X}|}{2}\log 2\pi-\frac{n}{2}\log|\Sigma_{\mathbf{X}}|-\frac{n}{2}\text{tr}((\Sigma_{\mathbf{X}})^{-1}\hat{\Sigma}_{\mathbf{X}}),$ (6)
[…]
where $n$ is the number of i.i.d. observations of $\mathbf{X}$, $\text{tr}(\dot{)}$ returns the trace of a matrix, and $\hat{\Sigma}_{\mathbf{X}}$ is the sample covariance estimated from data. We solve the optimization by gradient descent. We note that it is a non-convex optimization and thus we shall rely on multiple random initializations to approach the global optimum.
4.4 The Whole Procedure and Potential Usage
We will apply the proposed rank-based latent causal discovery method to SP100 stock price data, which suffers from the three demos - time resolution mismatch, non-stationarity, and the existence of latent variables.
To deal with time resolution, mismatch, we assume the underlying procedure follows Eq. 1 and we use the log return of each stock, formulated as $\text{return}_{t}=\log(\text{price}_{t}-\text{price}_{t-1})$. By doing so, our observation of return can be taken as linearly aggregated data, as in Eq. 2, and thus we can take our data as the i.i.d. case to apply instantaneous causal discovery methods.
Regarding the non-stationarity, we will apply Baysian change point detection Fearnhead (2006); Adams and MacKay (2007); Xuan and Murphy (2007) to identify the changing points in the time series data. Once we detect the change point, we partion the time series into segments, and thus within each segments the time series can be taken as stationary.
As for the existence of latent variable, we apply our rank-based latent causal discovery method, together with GIN condition to identify the direction. We further apply our propose method to estimate all the causal coefficients. Therefore, our final result include the causal graph, together with the coefficients of all the edges.
The potential usage of our result could be very broad. One significant usage would be estimating causal effects using only observational data, by making use of simulated interventions ((Pearl, 2009; Lopez de Prado, 2022)). This is a very promising direction and we recommend readers to refer to Lopez de Prado (2022) for how that can be achieved based on the knowledge of a causal model (which is the result we get) and observational distribution, especially in a financial context.
@@ 5 Experiments
In this section, we apply our proposed method to Sp100 stock price dataset and see how well we can address the three issues, i.e., time resolution, non-stationarity, and latent variables.
For the first issue, throughout the experiments we use the log daily/hourly return data, which can be written as the sum of the log minute return. Therefore, suppose the causal effects among stocks can be approximated as linear and happen instantaneously, according to Section 2, those instantaneous causal relations can still be recovered from the aggregated sum data. This can be verified by the reasonable results shown below. For the rest of this section, we focus on dealing with nonstationary relations and latent variables.
5.1 Results on Constraint-based causal Discovery from heterogeneous/NOnstationary Data (CD-NOD)
The scatter plot in Figure 1 highlights the potential for significant shifts in causal relationships among stocks during the Covid-19 pandemic. To further investigate this and explore what information can obtained in a nonstationary market context, in this subsection we focus on ten representative stocks from the SP100. These stocks are across various sectors that might be influenced by the pandemic:
  • Pharmaceuticals: Pfizer (PFE) and Merck (MRK). These companies are directly involved in the development and distribution of Covid-19 treatments and vaccines.
  • Entertainment and Streaming: Netflix (NFLX), Disney (DIS). The stay-at-home policies and social distancing measures increased demand for home entertainment.
  • Travel and Hospitality: Boeing (BA), Booking Holdings (BKNG). Disney (DIS) may also be counted here as for their resorts. These companies experienced significant downturns due to travel restrictions and reduced consumer travel interest.
  • Consumer Services and Retail: Walmart (WMT), Visa (V). These companies represent the consumer behavior and spending. Walmart saw changes in demand for essentials, while Visa saw overall consumer spending and economic activities.
  • Technology: Apple (AAPL). Represents the overall technology sector.
  • Energy: Chevron (CVX). May be impacted by the global oil demand and prices.
The CD-NOD analysis was conducted on these ten selected stocks, utilizing their log daily return data spanning from 2019 to 2023. The discovered causal graph and the estimated “driving forces” of changing causal modules are shown in Figure 3 and Figure 4, respectively.
Figure 3: CD-NOD discovered causal graph on the ten selected stocks from SP100 using data from 2019 to 2023. The top “DATE” is the date index serving as a surrogate for nonstationarity in the dataset. Variables that are direct child of “DATE” are those whose causal generating mechanisms (i.e., the conditional distribution of this variable give its other stock parental variables) are subject to changes over time.
Figure 4: Visualization of estimated driving forces of changing causal modules using CD-NOD’s Phase III. Each row indicates one stock whose causal module changes over time. The left panel shows the so-called “driving force”, i.e., the top-2 principle nonstationary components recovered by Kernel Nonstationary Visualization (KNV). Note that the lines primarily indicate “change points” in trends; they do not represent actual prices, nor do they carry a direct physical interpretation on the y-axis. The right panel shows the largest ten eigenvalues of the corresponding Gram matrix.
The discovered causal graph indicates four stocks whose causal generating mechanisms are subject to changes over time: Pfizer (PFE), Boeing (BA), Apple (AAPL), and Netflix (NFLX). The change on PFE may be because that the new drug (and especially Covid-19 vaccines) researches and approvals or health crises can suddenly alter the landscape. BA could be subject to changes due to the significant fluctuations in air travel throughout these recent years. NFLX saw the change in the rapidly evolving media streaming space, and this change may be especially significant during the Covid-19 times during to high demand. Similarly, AAPL operates in a highly dynamic tech market where where innovation and competition naturally shift over time.
To specifically demonstrate how the causal modules of these four variables change over time, CD-NOD’s Phase III is performed and the result is shown in Figure 4. Several of the identified changes align with known events and can be reasonably explained. For instance, the notable changes in Pfizer (PFE) during the early stages of 2020 and the middle of 2021 coincide with the emergence of COVID-19 and subsequent vaccine development milestones. The changes in Netflix (NFLX) (more specifically, the density $p$(NFLX|AAPL)) around late 2020 and early 2021 correspond to their increasing subscriber base. Other changes might not stem directly from the pandemic but could be influenced by various factors over time.
Across the whole causal graph, most of the edges (direct causal effects) seem also reasonable: The effect from PFE to BA, as echoed in Figure 1, may not typically be evident but make sense in the context of the Covid-19 pandemic, where the development of vaccines likely bolstered traveler confidence, impacting the aerospace sector. The direct relationships among Boeing (BA), Disney (DIS), and Booking Holdings (BKNG) may reveal the causal effects within the travel industry. Visa (V) being a joint effect of multiple stocks across different sectors demonstrates how consumer spending patterns, influenced by the broader economic environment, can reflect on payment processing companies. Some other edges lack a straightforward interpretation though, such as edges from Walmart (WMT) to Apple (AAPL) and Merck (MRK).
The limitations of the CD-NOD results presented here lie in that the analysis is restricted to just ten stocks, not the entire market. This constraint arises from the necessity to employ general, nonparametric conditional independence tests (e.g., the kernel-based conditional independence (KCI) test (Zhang et al., 2012)) to detect the nonlinear relationships associated with the surrogate variable (time index): while it’s reasonable to assume linearity among stock variables, the nonstationarity of causal modules, represented by the relationship of other variables on time index, tends to be inherently nonlinear. However, KCI-test has a cubic time complexity relative to sample size, which therefore limits the number of variables considered due to time constraints. To address this issue, we move forward from another perspective: with causal discovery that incorporates latent variables (note that aligned with the surrogate time index here, the nonstationarity can also be represented by a latent variable).
@@ 5.1 Results on Constraint-based causal Discovery from heterogeneous/NOnstationary Data (CD-NOD)
(a) Change point detection for SP100 from 2008 to 2013. The detected change point in 2008 can be explained as Global financial crisis and subprime mortgage crisis starting from 2007. The change point in 2009 can be explained as European debt crisis starting from 2009.
@@ 5.2 Baysian Change Point Detection
To utilize our rank-based latent causal discovery method we need to assume that our data is stationary; otherwise the nonstationary would induce rank deficiency constraints that cannot correctly reflect the underlying causal structure. Therefore, we first employ Baysian change point detection Fearnhead (2006); Adams and MacKay (2007); Xuan and Murphy (2007) to identify the changing points in the time series data. We take the average log daily return over all Sp100 stocks and take that as input to the Baysian change point detection.
@@
Our results are shown in Figure 5, where most of the change points detected can relate to real-life financial crises or events. Specifically, the change point in 2008 can be explained as Global financial crisis and subprime mortgage crisis starting from 2007. The change point in 2009 can be explained as European debt crisis starting from 2009. The detected change point in 2016 could relate to Chinese stock market crash in 2016. The change point in 2018 could relate to cryptocurrency meltdown. The change point in 2020 relates to Covid-19. The change point in 2022 could relate to the Russian-Ukraine war.
@@
The detected change points in 2018 and 2020 are significant and are well related to real-life events. Thus we split the SP100 data using these two change points and have three time series: (i) SP100 from Jan 2017 to Dec 2017, (ii) SP100 from Jan 2019 to Dec 2019, and (iii) SP100 from Jan 2021 to Dec 2021. We will refer to them as SP100-2017, SP100-2019, and SP100-2021 respectively in the following. As we do not detect any significant change point inside these three years, they can be considered as stationary time series.
@@
In the next section, we will show the causal discovery results using these three time series and compare the common part as well as the difference.
(a) Causal discovery result with causal coefficients for Sp100-2017.
(b) Causal discovery result with causal coefficients for Sp100-2019.
(c) Causal discovery result with causal coefficients for Sp100-2021.
@@
Figure 6: Discovered causal graph for Sp100-2017, Sp100-2019, and Sp100-2021, respectively. The three graphs have many things in common, and yet a big difference between 2019 and 2021 is that travel-related stocks are very causally related to pharmaceutical stocks (or pharm-related latent variables) after Covid-19, while before Covid-19 it is not the case.
5.3 Rank-based Latent Causal Discovery with Causal Coefficients
In this section, we apply our rank-based latent causal discovery method to the Sp100 time series data. In addition, we further apply the GIN condition to determine all the directions and estimate the coefficients of all the edges. For the change point detection we used log daily return but for rank-based latent causal discovery we use log houlry return such that we can have more datapoints within each time series. All the variables are normalized to have zero mean and unit variance.
Our first experiment is to show how the change point, e.g., Covid-19, influences the underlying graph. To this end we apply our method to the three time series data Sp100-2017, Sp100-2019, and Sp100-2021 respectively. To better compare the common parts and the difference, we only pick a subset of the Sp100 stocks. They are travel-related stocks (BA) colored in orange, pharmaceutical stocks (BMY, PFE, JNJ, MRK, GILD, ABBV, CVS, LLY, AMGN) colored in green, and info-tech stocks (ADBE, IBM, CSCO, NVDA, MSFT, AAPL, AMZN) colored in blue.
The result is shown in Figure 6. We can see that these three graphs have things in common. The most significant pattern is that all the stocks are basically clustered by the type of the company. For example, in Sp100-2017, there are basically two latent variables influencing all the pharmaceutical companies, and the same logic also applies to Sp100-2019 and Sp100-2021. The difference among the three graphs is even more interesting. One prominent difference is that for both Sp100-2017 and Sp100-2019, the BA, which is related to travelling, is more causally related to tech companies. Specifically, in 2017, BA is caused by AMZN, AAPL, and IBM and in 2019, BA is caused by L3, which is the parent of tech companies MSFT, IBM, and ADBE. In contrast, in 2021, BA is influenced by two latent variables, L2 and L4, as shown in (c) in Figure 6, where L2 and L4 are the parents of many pharmaceutical companies. This can be well explained by the change point caused by Covid-19. Once a new vaccination is approved or announced to be successful, the confidence in both the travel section and pharmaceutical industry increases at the same time; on the other hand, if there are setbacks in vaccine development or distribution, airline stocks may suffer due to anticipated continued travel restrictions.
Figure 7: Causal discovery result with causal coefficients for Sp100-2021, with stocks from five sections. They are pharmaceutical stocks (BMY, PFE, JNJ, MRK, GILD, ABBV, CVS, LLY, AMGN) colored in green, and info-tech stocks (ADBE, IBM, CSCO, NVDA, MSFT, AAPL, AMZN, INTC) colored in blue, energy companies (CVX, XOM) colored in purple, finance companies (COF, GS, AXP, BK, USB, BAC, JPM) colored in yellow, and consumer-related companies (TGT, COST, WMT, HD) colored in orange.
We further apply our method to Sp100-2021 with even more stocks from more sections. Specifically, they are pharmaceutical stocks (BMY, PFE, JNJ, MRK, GILD, ABBV, CVS, LLY, AMGN) colored in green, and info-tech stocks (ADBE, IBM, CSCO, NVDA, MSFT, AAPL, AMZN, INTC) colored in blue, energy companies (CVX, XOM) colored in purple, finance companies (COF, GS, AXP, BK, USB, BAC, JPM) colored in yellow, and consumer-related companies (TGT, COST, WMT, HD) colored in orange.
The discovered causal graph with all the causal coefficients are shown in Figure 7. As expected, all the stocks from the same section are well clustered. For example, all consumer goods fall under single Latent L6 with relatively high coefficients. All financial companies are under single Latent L3 with one sub group under another representative L4 which again is a child of L3. This also implies that BAC, USB, JPM share something in common that is not shared by other financial stocks. As for the energy companies XOM and CVX, we observe that they are highly related with a high coefficient 0.84. At the same time they are also causally related to some other companies including CSCO, JPM, and INTC, with relatively small strength of coefficient.
Based on our causal discovery result, we can unveil more interesting findings, such as estimating causal effect by simulated intervention, and validating results by further comparing to real interventions, as suggested in Lopez de Prado (2022). As it is beyond of the scope of this paper, we will leave it for future work.